OpenStudy (anonymous):

solve the equation by factoring 24x^2 -30x =0

OpenStudy (anonymous):

a. x=54,x=30 b. x=0 c. x=5/4, x=0 d. x=5/4

OpenStudy (anonymous):

24x² - 30x = 0 6x * (4x - 5) = 0 x = 0 or x = 5/4

OpenStudy (anonymous):

Since it's multiple choice, you could even cheat by plugging in the options they gave you for x and seeing if you get 0 or not.

OpenStudy (anonymous):

thanks